OpenStudy (anonymous):

Solve for n (5 2/3 + 8 1/4) + 9 3/4 = n + 18 @lizz123 @Jacob,*

OpenStudy (anonymous):

@KiaraIsABro

OpenStudy (anonymous):

23+ 2/3 = n + 18 23 + 2/3 -18 = n n = 5 and 2/3

OpenStudy (anonymous):

hope this helps!

OpenStudy (anonymous):

can you medal me if this is the correct answer?

OpenStudy (anonymous):

sorry yeah there you go

OpenStudy (anonymous):

Thanks! hope you get it right!

OpenStudy (anonymous):

@Jacob,* can you help me with another

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

8 3/4 + n + 2 3/8 + 4 12 = 23 1/8

OpenStudy (anonymous):

solve for n

OpenStudy (anonymous):

when you say + 4 12 what do you mean?

OpenStudy (anonymous):

4 1/2 sorry

OpenStudy (anonymous):

8 3/4 + n + 2 3/8 + 4 1/2 = 23 1/8

OpenStudy (anonymous):

i need it done in 8 mins at least before 10 mins

OpenStudy (anonymous):

and i have one other after that

OpenStudy (anonymous):

15 5/8 + n = 23 1/8 15 5/8 - 15 5/8 + n = 23 1/8 - 15 5/8 n = 7 1/2

OpenStudy (anonymous):